OpenStudy (anonymous):

A rectangular field bounded on one side by a lake is to be fenced on 3 sides by 800 meters of fence. What dimensions will produce a maximum area?

OpenStudy (anonymous):

x + 2y = 800 m ( of fence) Area A = x * y -> A = ( 800 - 2y) * y = 800y - 2y² Max Area when A' = 0 @Unionjk Now it's so simple to solve for y, x dimensions :)